Thanks for your help in advance!

Describe the sampling distribution of p. Assume the size of the population is 20,000 n 800, p 0.1 Choose the phrase that best describes the shape of the sampling distribution of p below. ns0.05N and np(1 p)10. 0.05N and np(1-p)210 O A. Not normal because B. Approximately normal because n s 0.05N and np(1-p) < 10. C. Not normal because n (YD. Approximately normal because n 0.05N and np(1-p)210. Determine the mean of the sampling distribution of p. μ,-0.1 (Round to one decimal place as needed.) Determine the standard deviation of the sampling distribution of p On-U (Round to three decimal places as needed.)
Answer #1

Given sample n=800 & p=0.1

Mean of the sampling distribution of p^ is= p=0.1

Standard deviation of the sampling distribution of p^ is

= sqrt ( p ( 1-p)/n)

=sqrt (0.1×0.9/800)

=Sqrt (0.00011)

=0.0106 { \simeq } 0.011